OpenStudy (lexilove13):

Algebra 2 Help Please! Binomial Distribution.

OpenStudy (lexilove13):

According to a national survey, 16% of U.S. adults use the internet to make telephone calls. In a random survey of 100 adults what is the probability that at most 20 of them use the internet to make telephone calls?

OpenStudy (lexilove13):

@jim_thompson5910 @Miracrown

jimthompson5910 (jim_thompson5910):

do you have a calculator like a TI83 or TI84?

OpenStudy (lexilove13):

N=100 P=0.16 Np=100*0.16 = 16 > 5 N(1-p)=100*0.84=84 > 5 Mean u = np = 16 Standard deviation o = sqrt (np(1-p)) = sqrt(100(0.16)(0.84)) = 4 Z= 20-16/4 = 4/4 = 1 P(x<_20) = *Not sure how to answer this part*

jimthompson5910 (jim_thompson5910):

oh that works too

OpenStudy (lexilove13):

No I do not! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me look over your work

jimthompson5910 (jim_thompson5910):

sqrt(100(0.16)(0.84)) is not equal to 4

OpenStudy (lexilove13):

It was about equal to but the sign would not work

jimthompson5910 (jim_thompson5910):

sqrt(100(0.16)(0.84)) = 3.67 roughly

jimthompson5910 (jim_thompson5910):

since we are approximating the binomial distribution with a normal distribution, we need to use the continuity correction factor https://people.richland.edu/james/lecture/m170/ch07-bin.html

jimthompson5910 (jim_thompson5910):

Using the continuity correction factor, we get \[\Large P(X \le 20) \to P(X < 20+0.5)\] \[\Large P(X \le 20) \to P(X < 20.5)\]

jimthompson5910 (jim_thompson5910):

x = 20.5 mu = 16 sigma = 3.67 z = (x - mu)/sigma z = (20.5 - 16)/3.67 z = 1.23 P( X < 20.5) = P(Z < 1.23) You will then use a table like this one https://www.stat.tamu.edu/~lzhou/stat302/standardnormaltable.pdf to compute P(Z < 1.23) according to that table, P(Z < 1.23) is approximately 0.89065 So P(X < 20.5) = 0.89065 approximately